OpenStudy (anonymous):

in 1930 the average cost for a gallon of gasoline was $0.25. in 2007 the average cost for a gallon of gasoline is $2.60 what is the percent of increase rounded to the nearest whole number

OpenStudy (tkhunny):

0.50 / 0.25 = 2 = 100% Increase 1.00 / 0.25 = 4 = 300% Increase 2.00 / 0.25 = 8 = 700% Increase Seeing a pattern?

OpenStudy (anonymous):

geometric...